OpenStudy (anonymous):

The sum of two number is 36. Four times the smaller is 1 less than the larger. What are the number?

OpenStudy (unklerhaukus):

\[a+b=36\]\[4a=b-1\] \[4a+1=b\]\[a+4a+1=36\]\[(1+4)a=36-1\]\[a=\frac {36-1}{1+4}\]\[b=36-a\]